two rectangles have the same unknown area
but the length of one is (6x+6)ft and width 4ft and the other has length 9ft and width 4xft. How do I fine what the variable x is?

Answer by jsmallt9(3758)

The area of the first rectangle would be (6x+6)*4 = 24x+24.
The area of the second triangle would be 9*4x = 36x
We've been told that the two rectangles have the same area. So:
24x+24 = 36x

Now we just solve for x. Subtracting 24x from each side:
24 = 12x
Dividing by 12:
2 = x